Grade 9Physical Chemistry

a quantity of gas is collected in a graduated tube over the mercury. The volume of a gas at 20°C is 50ml and the level of mercury is 100mm above the outside mercury level. The atmospheric pressure is 750mm . volume at STP is????(given-R=0.083 atm/K/atm )

Sridhar Karumanchi
9 Years agoGrade 9

1 Answer

Sridhar Karumanchi
9 Years ago

P1 = 650 
V1 = 50 ml 
T1 = 20 C + 273 = 293 K 
P2 = 760 (convert)
V2 = ? 
T2 = 273 K 
Ideal gas law: 
P1 V1 / T1 = P2 V2 / T2 

V2 = 650 * 50 *273 /293 *760
V2 = 39.84 ml